OpenStudy (thatgirl21):

Please help, will give medals and fan. -there are multiple parts to this question. If someone could just show me what to do on one of them that would help so much. If f(x)=6x-3 and g(x)=11-2x and h(x)=5x^2 a.) (f*g)(x)=? b.) (f*h)(2)=? c.) (g*h)(x)=? d.) (g*f*h)(x) e.) (g*f)(-2) f.) (g*g)(x)

satellite73 (satellite73):

is it \[f\times g(x)\] or is it \[f\circ g(x)\] make s a big difference

OpenStudy (thatgirl21):

it has a circle

OpenStudy (thatgirl21):

like the second one you typed

satellite73 (satellite73):

that is what i though

satellite73 (satellite73):

lets do this one \[(f\circ g)(x)\]

satellite73 (satellite73):

the first step is to get it out the the circle notation, and write down what it really means \[f\circ g(x)=f(g(x))\]

satellite73 (satellite73):

that is always the first step, no matter what \(f\) and \(g\) are the second step is to replace the general \(g(x)\) by the spefic one you have, namely \(11-2x\) so \[f(g(x))=f(11-2x)\]

satellite73 (satellite73):

the third and only abstract step is to replace the \(x\) in \(f(x)\) by \(11-2x\) think \[f(\spadesuit)=6\spadesuit-3\] so \[f(11-2x)=6(11-2x)-3\]

satellite73 (satellite73):

last step is to clean it up with algebra distribute combine like terms i'll leave that to you

OpenStudy (thatgirl21):

63-12x?

OpenStudy (thatgirl21):

So for #2. You do f(g(2)) which is f(20)? Is that how you start this one? and then 6(20)-3 which is 117

OpenStudy (thatgirl21):

Not sure if I did that right

satellite73 (satellite73):

yes, \(63-12x\) is right

satellite73 (satellite73):

oops it is not \[f(g(2))\] is it? it is \[f(h(2))\]

OpenStudy (thatgirl21):

yes! f(h(2))

satellite73 (satellite73):

your answer is right

OpenStudy (thatgirl21):

Awesome, thank you so much. For c g(h(x)) I did 11-2(5x^2) and got -10x^2+11 and that looks correct to me. I'm a little confused about d though

satellite73 (satellite73):

C is right

satellite73 (satellite73):

D you just got to do it twice \[f(g(h(x))\] is the first step then replace \(h\) by \(5x^2\) then ...

satellite73 (satellite73):

oops i had them in the wrong order it is \[g(f(h(x))\] but that doesn't change the first step \[g(f(5x^2))\]

satellite73 (satellite73):

if you want i will check it

OpenStudy (thatgirl21):

I got 60x^2+17, not sure if I did that right though /: I did 11-2(6(5x^2)-3) multiplied 6 by 5x^2 and ended up with 11-2(30x^2-3) then multiplied 2 by (30x^2-3) and combined like terms to get 60x^2+17

satellite73 (satellite73):

probably right i would do it one step at a time, may be easier

satellite73 (satellite73):

\[f(h(x))=f(5x^2)=6\times 5x^2-3=30x^2-3\]

satellite73 (satellite73):

then \[g(30x^2-3)=11-2(30x^2-3)\] think you are off by a minus sign

satellite73 (satellite73):

otherwise good, maybe you just dropped it by mistake \[-60x^2+17\] is what i got

OpenStudy (thatgirl21):

oh I wrote my minus sign down but forgot about sticking with it at the very end!

satellite73 (satellite73):

ok well it looks like you got this pretty quickly

OpenStudy (thatgirl21):

Thank you so much for making this easier (: one more question if you don't mind, for the last one do I just do 11-2(11-2x)?

satellite73 (satellite73):

yes

OpenStudy (thatgirl21):

Okay awesome! I really appreciate it.

satellite73 (satellite73):

my pleasure, glad to help someone who helps themselves